2. Calculatingand categorizing inflation This table indicates the historical level of the Consumer Price Index (CPI) for theWhat rates of inflation for 1924 would be consistent with hyperinflation? Check all that apply. O 15.0% 120.0% 0 -1.8% O 100.
Answer #1

Inflation Rate Change in Price Level Year 1921 1922 1923 CPI 17.9 16.8 17.1 -6.1% Deflation 1.8% Inflation What rates of infl

inflation rate between 1921 and 1922 = (16.8 - 17.9)/17.9 = -6.1%

inflation rate between 1922 and 1923 = (17.1 - 16.8)/16.8 =  1.8%

Hence, inflation rate in 1924, to be less than 1.8% and greater than 0%. will create disinflation between 1923 and 1924.

What rates of inflation for 1924 would be consistent with hyperinflation? Check all that apply. 15.0% 120.0% -1.8% 100.0%